# On the Salient Limitations of the Methods of Assembly Theory and their Classification of Molecular Biosignatures

### v8

@@ A.1 Mischaracterisations
A.2 Expected false positives from complexity-deceiving molecules with arbitrarily high statistical significance
Indeed, as we move beyond the realm of pure stochastic processes, complexity distortions become even more problematic [zenil_kiani_2017, Abrahao2021publishednat, Abrahao2021darxiv, Abrahao2020cAIDistortionsCN]. As demonstrated in Theorem A.4 and Corollary A.5, there are (sufficiently large) deceiving molecules the complexities of whose respective generative processes arbitrarily diverge from the assembly index that the assembly pathway method assigns to them. We demonstrate in Section A.2.1 that there are (sufficiently large) deceiving molecules the complexities of whose respective generative processes arbitrarily diverge from the assembly index that the assembly pathway method assigns to them. By a generative process [Abrahao2021bpublished] we mean any process that can be implemented, computed, or emulated by another equivalent (or identical) process so that it generates the pathway assembly and its object. In accordance with the assumptions and rationale in [cronin, marshall_murray_cronin_2017, Marshall2019], those generative processes are exactly those composed of the assembling processes allowed (i.e., deemed physically possible by assembly theory’s chosen method) so that they in the end result in the constitution of molecules.
Notice that our results in Section A.2.1 hold for any of such processes that belong to (abstract or physically implemented) computational classes much weaker than that of Turing machines, e.g. resource-bounded Turing machine [Zenil2020, Allender2011, soler, Calude2002main, Burgin2009, liandvitanyi] or simple forms of finite automata [Wolfram2002, Hutchison2006]. Nevertheless, if the class of generative processes is constituted by processes that are also capable of universal computation, then Corollary A.5 shows the complexity distortions (and, therefore, also the deceiving phenomena we explain below) can be equally bad or even worse. Therefore, AT fails to capture, both in theory (when resources can be unbounded) and in practice (when resources are limited), the minimality that is necessary for a complexity measure that may be claimed to be unambiguous and observer-independent.
While the calculation of MA may be prone to false negatives—due to partial fragmentation in energy collision analysis and the restriction to counting only valence rules in molecule synthesis (ignoring other chemical conditions)—this does not pose a challenge to the central claim made in [cronin]. Instead, MA aims at avoiding underestimation of the amount of molecules that result from random or abiotic processes, hence aiming at avoiding false positives. Thus, by directly tackling their central claims and this motivation, one demonstrates in Theorem A.4 that MA in general fails to avoid false positives in the specific sense that it may not be able to distinguish a “complex” object that is in fact the outcome of randomly generated (resource-bounded) generative processes.
Under the same assumptions as in [cronin, Marshall2019], we construct in Theorem A.4 a deceiving molecule that has a much larger MA value in comparison to the minimal information sufficient for a randomly generated generative process to single-handedly construct this molecule. Whatever arbitrarily chosen method is used to calculate the statistical significance level, the MA of this molecule is large enough to make the expected frequency of occurrence (estimated via the arbitrarily chosen AT) diverge from the actual probability. In this case, AT would consider such a molecule “biotic”, resulting from extrinsic factors that increase biases toward certain pathways or that constrain the range of possible joining operations, although its sole underlying generative process in fact results from fair-coin-toss random events. This proven existence of false positives due to such a deceiving phenomenon is corroborated by our empirical findings, which show that MA displays a behaviour that is both structurally and empirically similar to traditional statistical compression methods. Indeed, the latter methods are already known to present distorted values [zenil_kiani_2017], performing worse than more recent algorithmic-based methods [Zenil2020]. Thus, they are prone to overestimating complexity, and consequently to presenting false positives for “high”-complexity objects which are in fact simple.
Nevertheless, note that it is true that there are computable (lossless) encodings of a source, such as Huffman coding, that are proven to be optimal on average, but only if one knows a priori that the underlying processes generating the objects are purely stochastic (in particular, when one knows beforehand that the conditions of the source coding theorem are satisfied [Cover2005]). In this case, one can show that the minimum expected size of the encoded object converges to its expected algorithmic complexity [Calude2002main, Cover2005]. However, pure stochasticity is too strong an assumption, or does not realistically represent the generative processes of molecules.
On the one hand, living systems are complex systems consisting of multiscale, multi-nested processes that are unlikely to be reducible to simplistic and intrinsic statistical properties such as those suggested by AT. One cannot conceive a measure that only looks at the internal structure of an agent isolated from its environment and how it interacts with its external medium to determine its (non)living nature. Especially in the context of complex systems like living organisms, organic molecules may be the byproduct of intricate combinations or intertwinements of both deterministic/computable and stochastic processes that govern the behaviour of the entire organism and its environmental surroundings [Walker2012nat, bdm].
On the other hand, in case a sufficiently complex (e.g., at the level of those we demonstrate in Section A.2.1) environmental catalytic condition plays the role of the above extrinsinc factor (which increases the bias toward the construction of a more complex molecule), this level of complexity would be completely missed by the capabilities of simplistic measures such as MA, thereby rendering it as prone to false positives. More careful and deeper arguments regarding simplicity, recursivity, and the emergence of modularity in life have been advanced and are better grounded in a theoretical and methodological framework advanced in [santiago], where it was shown that exploiting first principles of computability and complexity theories, modular properties in evolutionary systems may be explained.
As shown by Corollary A.5, the deceiving phenomena can be equally bad or even worse in case the molecules are byproducts of complex systems that are somehow capable of universal computation. For example, in the case of advanced civilisations that are capable of artificially constructing living beings by computable processes, simplistic complexity measures such as MA can be intentionally misled with respect to what actually should be measured; or in the case of extrinsic environmental catalytic processes whose chemical dynamics’ complexity are comparable to those of some randomly generated finite automata [Wolfram2002].
Our also results suggest the presence in more realistic resource-limited scenarios of a property known to occur in theoretical complex systems science [Abrahao2021bpublished]: in the context of generative processes that are not purely stochastic, and are capable of displaying complexity or computation capabilities at the same level of those of the observers, there is no such thing as a generally optimal complexity measure that cannot be improved upon, since computable complexity measures are dependent on the observer (or the chosen formal theory). For example, without the necessary conditions being satisfied by the underlying stochastic process, one cannot generally guarantee such a convergence between the expected size of the encoded object and the expected algorithmic complexity that is assured by the source coding theorem.
A.2.1 Mathematical framework and assumptions
Nevertheless, in an ideal (resource-unbounded) case in which the whole algorithm space indeed constitutes the set of all possible generative processes for constructing the assembly space (e.g., when chemical processes are able to achieve the capability of effecting universal computation in the real world [Walker2012nat, Wolfram2002]), we also show in Corollary A.5 that the deceiving phenomenon hold in the same way (or can even be worse).
A deceiving phenomenon akin to the one employed in Theorem A.4 can be found in [Abrahao2021darxiv] based upon the same AIT principles in [zenil_kiani_2017], where sufficiently large datasets were constructed so that they deceive statistical machine learning methods into being able to find an optimal solution that in any event is considered global by the learning method of interest, although this optimal solution is in fact a simpler local optimum from which the more complex actual global optimum is unpredictable and diverges.
This phenomenon is also related to the optimality of the algorithmic complexity as an information content measure that takes into account the entire discrete space of computable measures [Downey2010, Chaitin2004], or the maximality of the algorithmic probability as a probability semimeasure on the infinite discrete space of computably constructible objects, as demonstrated by the algorithmic coding theorem [liandvitanyi, Downey2010, Calude2002main].
In order to achieve our results, we base our theorems on mathematical conditions that are consistent with the assumptions and results in [cronin, Marshall2019]. The first assumption that we specify with the purpose of studying a worst-case scenario is that the assembly space should be large enough so as to include those molecules (or objects) with sufficiently large MA (along with its associated sufficiently low pathway probability of spontaneous formation) relative to the algorithmic complexity of the deceiving program. For the sake of simplicity, we assume that the nested family $\mathcal{S}$ of all possible finite assembly spaces from the same basis (i.e., the root vertex that represents the set of all basic building blocks) is infinite computably enumerable. However, an alternative proof can be achieved just with the former—and more general—assumption that the assembly space may be finite but only needs to be sufficiently large in comparison to the deceiving program. Indeed, our assumption is in consonance with the authors’ motivation (and/or assumption) that “biochemical systems appear to be able to generate almost infinite complexity because they have information decoding and encoding processes that drive networks of complex reactions to impose the numerous, highly specific constraints needed to ensure reliable synthesis” [cronin].
Thirdly, in accordance with the arguments in [cronin, marshall_murray_cronin_2017, Marshall2019] that the computability and feasibility of their methods is an actual advantage in comparison with other complexity measures, here we likewise adopt these same assumptions so that the following are computable procedures:
  • deciding whether or not a finite assembly space (or subspace) is well formed according to the joining operation rules that are allowed to happen;111 For example, one can employ the same criteria and allow the same rules established in [cronin].
  • calculating MA of a finite molecule (i.e., a finite object) in a well-formed assembly space (or subspace);222 For example, as defined in [Marshall2019, Definition 19].
  • calculating the chosen approximation of MA (e.g., the split-branch version) of a finite molecule in a well-formed assembly space (or subspace);333 For example, as in [cronin, SI] and [Marshall2019, Section 4.2].
  • calculating an upper bound for the pathway probability of spontaneous formation of a molecule in the denumerable nested family of possible finite assembly spaces;444 For example, this can be done by employing the methods developed in [Marshall2019, SI] and [cronin]
  • calculating the significance level for a frequency of occurrence of a molecule in a sample so that this empirical probability distribution (i.e., the type of the sample) diverges from the pathway probability distribution of spontaneous formation of the molecule;555 For example, by using a maximum-likelihood method or using the probability that a sample occurs with KL divergence larger than $\epsilon$ [Cover2005, Goodfellow2016].
A.2.2 Definitions and notation
Besides the notation from [Marshall2019] for assembly theory, we also employ the usual notation for Turing machines and algorithmic complexity.
Respectively, as in [Marshall2019, Definition 11] and [Marshall2019, Definition 15], let $\left(\Gamma,\phi\right)$ denote either an assembly space or assembly subspace. From [Marshall2019, Definition 19], we have that $c_{\Gamma}\left(x\right)$ denotes the assembly index of the object $x$ in the assembly space $\Gamma$.
Note that assembly spaces are finite. So, from our assumptions, we need to define a pathway assembly that can deal with arbitrarily large objects. To this end, let $\mathcal{S}=\left(\mathbf{\Gamma},\mathbf{\Phi},\mathcal{F}\right)$ be an infinite assembly space, where every assembly space $\Gamma\in\mathbf{\Gamma}$ is finite, $\mathbf{\Phi}$ is the set of the correspondent edge-labeling maps $\phi_{\Gamma}$ of each $\Gamma$, and $\mathcal{F}=\left(f_{1},\dots,f_{n},\dots\right)$ is the infinite sequence of embeddings [Hodges1993] (in which each embedding is also an assembly map as in [Marshall2019, Definition 17]) that ends up generating $\mathcal{S}$. That is, each $\,f_{i}\colon\left\{\Gamma_{i}\right\}\subseteq\mathbf{\Gamma}\to\left\{\Gamma_{i+1}\right\}\subseteq\mathbf{\Gamma}\,$ is a particular type of assembly map that embeds a single assembly subspace into a larger assembly subspace so that the resulting sequence of nested assembly subspaces defines a total order $\preceq_{\mathcal{S}}$, where
[…]
Let ${\Gamma^{*}}_{x}$ denote a minimum rooted assembly subspace of $\Gamma$ from which the assembly index $c_{\Gamma}\left(x\right)$ calculates the augmented cardinality and that its longest rooted paths $\gamma_{x}$ ends in the arbitrary object $x\in V\left(\Gamma\right)$ as in [Marshall2019, Definition 19].
[…]
We have that the (prefix) algorithmic complexity, denoted by $\mathbf{K}\left(x\right)$, is the length of the shortest prefix-free (or self-delimiting) program $x^{*}\in\mathbf{L}$ that outputs the encoded object $x$ in a universal prefix Turing machine $\mathbf{U}$, i.e., $\mathbf{U}\left(x^{*}\right)=x$ and the length $\left|x^{*}\right|=\mathbf{K}\left(x\right)$ of program $x^{*}$ is minimum. In addition, the algorithmic coding theorem [Downey2010, Chaitin2004, Calude2002main, liandvitanyi] guarantees that
If the language $\mathbf{L}^{\prime}$ is a proper subset of $\mathbf{L}$ such that language $\mathbf{L}^{\prime}$ running on $\mathbf{U}$ is not able to decide every problem of Turing degree $\mathbf{0}$, then we have that $\mathbf{L}^{\prime}$ is not a universal programming language, and the machine $\mathbf{U}$ defined upon language $\mathbf{L}^{\prime}$ is a Turing submachine $\mathbf{U}/f$ [Abrahao2016nat], where $f$ is the partial function that computes the function $\mathbf{U}\left(x\right)$ for $x\in\mathbf{L}^{\prime}$ as input. In other words, a Turing submachine is a Turing machine that can receive inputs in its tape (and possibly simulate other machines), but it is not universal. Weaker than Turing degree $\mathbf{0}$, a submachine can only compute problems in a subrecursive class of problems [Abrahao2016nat]. Thus, note that resource-bounded machines or total Turing machines are particular cases of Turing submachines [Abrahao2016nat].
As a consequence of the above definitions, we define the (prefix) sub-algorithmic complexity666 See also [Abrahao2016nat] where this terminology is also employed. $\mathbf{K}_{f}\left(x\right)$ to be the length of the shortest prefix-free (or self-delimiting) program $x^{*}\in\mathbf{L}^{\prime}$ that outputs the encoded object $x$ when run on the Turing submachine $\mathbf{U}/f$ (i.e., $\mathbf{U}/f\left(x^{*}\right)=x$ and $\left|x^{*}\right|_{f}=\mathbf{K}_{f}\left(x\right)=\min\left\{\left|w\right|\,\middle|\,\mathbf{U}\left(w\right)=x,\,w\in\mathbf{L}^{\prime}\right\}$).777 If there is no program in $\mathbf{L}^{\prime}$ that can output an object $x$, then one defines $\mathbf{K}_{f}\left(x\right)=\infty$. Thus, note that resource-bounded variants of the algorithmic complexity [Calude2002main, Burgin2009, liandvitanyi] are particular cases of sub-algorithmic complexity. Analogously, we will have that the sub-universal (a priori probability) distribution upon language $\mathbf{L}^{\prime}$ is defined on the sub-universal a priori probability for each value $x$, which are given by
$\sum\limits_{\small\begin{array}[]{c}\mathbf{U}\left(p\right)=x;\\ p\in\mathbf{L}^{\prime}\end{array}}C\,\frac{1}{2^{\left|p\right|}}\text{ ,}$ (2)
where $\mathbf{L}^{\prime}\subseteq\mathbf{L}$ and $C\geq 1$ is a normalizing constant as in [Abrahao2017publishednat, Definition 3.6, Section 3.2.1] to ensure it is a probability measure and not a probability semimeasure. Note that, if $\mathbf{L}^{\prime}=\mathbf{L}$, then one obtains the usual universal distribution instead of its subrecursive version.
Following these notions, we can now define the submachine that compute the allowed (physical, chemical, and/or biological) generative processes of an assembly space that assembly theory may arbitrarily choose. Following the same rationale from [cronin, marshall_murray_cronin_2017, Marshall2019], those generative processes are exactly composed of the processes allowed by our current knowledge of physics so that they in the end result in the constitution of molecules.
As assumed to the above, and in accordance with the claims in [cronin, marshall_murray_cronin_2017, Marshall2019], notice that the computability and feasibility of assembly theory’s methods directly implies the existence of such a submachine. Given any set of (physical, chemical, and/or biological) rules for assembling molecules that assembly theory arbitrarily chooses to be based on and that is consistent with our current knowledge of physics, one has it that for any particular pathway assembly resulting in a molecule, there is a correspondent program running on a submachine $\mathbf{U}_{\Gamma}$ for which one can apply the methods described in [cronin, marshall_murray_cronin_2017, Marshall2019] to calculate the pathway probability and so on. This existence is trivially guaranteed to hold as long as the calculation of the assembly index is computable given an assembly space, which is already an assumed condition in [cronin, marshall_murray_cronin_2017, Marshall2019]; moreover, any upper bound for the computational resources necessary to compute the assembly index straightforwardly implies the existence of an upper bound for the computational resources necessary for a program to run on $\mathbf{U}_{\Gamma}$ (i.e., submachine $\mathbf{U}_{\Gamma}$ is a type of resource-bounded Turing machine).
For the sake of simplifying notation, $\mathbf{U}_{\Gamma}$ denotes the Turing submachine $\mathbf{U}/{f}_{\Gamma}$. In this case, the function $f_{\Gamma}$ is the partial function that returns what $\mathbf{U}$ can compute with some $x\in\mathbf{L}_{\Gamma}$ as input, where $\mathbf{L}_{\Gamma}\subseteq\mathbf{L}$ is a (non-universal) programming language such that every allowed generative process—i.e., every generative process that is deemed physically possible by assembly theory’s chosen method, in case the objects are molecules—of an assembly space is bijectively computed (or emulated) by a corresponding $\mathbf{U}\left(x\right)$. In other words, for every allowed generative process that can assemble objects into building another object, there is a program $x\in\mathbf{L}_{\Gamma}$ that computes (or emulates) this process. Conversely, for every $x\in\mathbf{L}_{\Gamma}$, one also has it that there is a corresponding generative process allowed by assembly theory, process which is computed (or emulated) by program $x\in\mathbf{L}_{\Gamma}$. Notice that in the forthcoming Theorem A.4, we will show that if a certain generative process is equivalent (in terms of complexity and computation capability) to a particular type of algorithm running on this resource-bounded submachine, then the deceiving phenomenon takes places. Although our theoretical results prove the existence of false positives in the case this condition is met, our empirical results shown in the main article not only corroborate this prediction but in fact suggest this theoretical finding (or some weaker version of it) may be happening in much lower computational classes. This occurs because the assembly index either underperforms or does not outdistance other measures of a statistical nature, where distortions in complexity estimations are already well known to occur—which in particular is one of the reasons complexity science moved on from simplistic statistical measures and remains an evolving field of research.
A.2.3 Theoretical results
Lemma A.1.
Proof.
holds. In order to demonstrate that $p$ always halts, just note that $\mathcal{S}$ is infinite computably enumerable. Also, for any value of $c_{\Gamma^{\prime}}\left(z\right)$ for some $z\in V\left(\Gamma^{\prime}\right)\subset V\left(\mathcal{S}\right)$, there is only a finite number of minimum rooted assembly subspaces (starting on any object in $B_{\mathcal{S}}$ and ending on $z$) whose augmented cardinality is $c_{\Gamma^{\prime}}\left(z\right)$, where $B_{\mathcal{S}}$ is the basis (i.e., the finite set of basic building blocks [cronin]) of $\mathcal{S}$. This implies that there is an infinite number of distinct values of $c_{\Gamma^{\prime}}\left(z\right)$. Now, let $\mathrm{p}_{y}=\left<k,\mathbf{F},p\right>$. Finally, from Equation 4 and basic properties in AIT, we have it that
holds for some sufficiently large $k$. ∎
Lemma A.2.
Let the conditions for Lemma A.1 be satisfied. Let
be an arbitrary probability measure on the set of pathways in $\mathcal{S}$ and $\mathrm{p}_{\mathbf{P}}$ a program that computes a computable function that outputs an upper bound for $\mathbf{P}$ such that for every $\epsilon^{\prime}>0$ and $\Gamma\subset\mathcal{S}$, there are $\Gamma^{\prime}\subset\mathcal{S}$ and $x\in V\left(\Gamma^{\prime}\right)$ with $\Gamma\subseteq\Gamma^{\prime}$ and $\mathbf{P}\left({\Gamma^{*}}_{x}\right)\leq\mathbf{U}\left(\left<{\Gamma^{*}}_{x},\mathrm{p}_{\mathbf{P}}\right>\right)<\epsilon^{\prime}$. Let $01\geq\epsilon>$ 0 be an arbitrary encodable real number. Let $k\in\mathbb{N}$ be an arbitrarily large natural number. Then, there are a program $\mathrm{p}_{\epsilon}$, $\Gamma\subset\mathcal{S}$, and $y\in V\left(\Gamma\right)$ such that Lemma A.1 is satisfied with $y$ and
and
Proof.
Let $p^{\prime}$ be a bit string that represents an algorithm running on a prefix universal Turing machine $\mathbf{U}$ that receives $\mathrm{p}_{\mathbf{P}}$, $\epsilon$ and $k$ as inputs. Then, it enumerates the assembly pathways ${\Gamma^{*}}_{x}$ in $\mathcal{S}$ such that $\left|p\right|+\left|\mathbf{F}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+\left|p^{\prime}\right|+\left|\mathrm{p}_{\mathbf{P}}\right|+\mathbf{O}\left(\log_{2}\left(\epsilon\right)\right)+k+\mathbf{O}(1)\leq c_{\Gamma}\left(x\right)$, $\mathbf{U}\left(\left<{\Gamma^{*}}_{x},\mathrm{p}_{\mathbf{P}}\right>\right)<\epsilon$, and Lemma A.1 holds for $x$ given $k$. Finally, it returns this first ${\Gamma^{*}}_{x}$ in this enumeration. Now, let $\mathrm{p}_{\epsilon}=\left<\mathrm{p}_{\mathbf{P}},\epsilon,k,p^{\prime}\right>$. Therefore, in addition to the arguments in the proof of Lemma A.1, the desired theorem follows from the fact that program $p^{\prime}$ always halts because of our initial assumptions on program $\mathrm{p}_{\mathbf{P}}$ and the probability distribution given by $\mathbf{P}$. ∎
Lemma A.3.
Let the conditions for Lemmas A.1 and A.2 be satisfied. Let $\mathbf{F}^{\prime}\supseteq\mathbf{F}$ be a formal theory that also includes the chosen statistical method, the criteria for the arbitrarily chosen statistical significance level, the chosen computable method for approximating $\mathbf{P}$ from above with program $\mathrm{p}_{\mathbf{P}}$, and the acceptable maximum error $\mathscr{E}\in\mathbb{N}$ for an overestimation of the complexity of an object in $\mathcal{S}$. Then, there are a program $\mathrm{p}_{d}$, $\Gamma\subset\mathcal{S}$, and $y\in V\left(\Gamma\right)$ such that Lemma A.2 is satisfied with $y$ and $\mathbf{F}^{\prime}$ decides that the divergence of the (sub-)universal distribution from $\mathbf{P}$ is statistically significant, where $\mathbf{U}\left(\mathrm{p}_{d}\right)={\Gamma^{*}}_{y}$ and $\left|\mathrm{p}_{d}\right|+\mathscr{E}<c_{\Gamma}\left(y\right)$ hold.
Proof.
[…]

### v10

@@ Appendix A Mischaracterisations
In [18], the authors of Assembly Theory attempted to address what they thought were some concerns about an early preprint version of this manuscript. According to the authors, the statistical tools used were not appropriate, and the figures and comparisons above confound different types and sources of data with their application being unique in that it can be applied to data closer to what they think is the physical process, this is not correct. Regarding the statistical tools, even conceding (which we are not), that other statistical comparison methods are possible, it is not possible to manufacture desired results with one tool that with another one is found to be completely different. The authors seem to suggest that the greatest value of this paper is that some measures outperformed their Molecular Assembly (MA) index, and not that every other statistical index including some introduced in the 60s, applied to any chemical input data (not only mass spectra), reproduces and replicates what their paper [32] introduced as its main contribution.
Appendix B On dictionary-based algorithms
The assembly method derived from the ‘Assembly Theory’ proposed by the original authors [34] consists roughly in finding a pattern-matching generative grammar behind a string by traversing and counting the number of copies needed to generate its modular redundancies, decomposing it into the statistically smallest collection of components that reproduce it without loss of information by finding repetitions that reproduce the object from its compressed form.
For purposes of illustration, let us take the example of ABRACADABRA, which the original authors have also used [34]. For molecular assembly (MA) to succeed it needs to have a discriminator and classifier able to characterise each repetition of $A$ and $N$ as the same, where $N$ is another character or some sub-unit of the structure with the same frequency as $A$ (e.g., a two-letter unit containing $A$, such as $AB$ or $RA$). In the ABRACADABRA example, MA deconstructs the sequence into unique blocks of five possible characters by adding a new character in subsequent steps, such that the minimal number of steps, considering only the frequency of the largest repeated block size (ABRA) just as it is done for LZ compression algorithms. The repeated binary or tertiary recursive structures (i.e., blocks of 2 or 3 letters) within the sequence, such as AB, RA, or BRA, are ignored in MA’s minimal path search as in LZ.
The proposed Molecular Assembly (MA) [32], the assembly index, and the Assembly Theory in general [34] fall into the category of dictionary-based entropy encoding indexes and approaches and are indistinguishable from an implementation motivated by and based upon the principles of algorithmic complexity using LZ compression. In Section B, we have shown the behavioural similarity between statistical compression and MA results consistent with the theoretical findings [7]. These popular statistical algorithms have been universally used for data compression and as computable estimations to algorithmic complexity [30] and logical depth [51, 41] including for classification purposes of living systems [16, 46, 51] of which AT is a special and, mathematically speaking, weak estimation.
Figure 3: ABRACADABRA tree diagrams for AT (A) and dynamic Huffman coding (B), both computable measures trivial to calculate. Huffman’s was the first dictionary-based coding algorithm and is an optimal coding method able to characterise every statistical redundancy, including modularity, independent of such copy data representation [7].The (molecular) assembly index has been proven to be equivalent to LZ77/LZ78 [7]. In this example, Huffman’s (which is also a sequential lossless compression algorithm that traverses strings from left to right) collapses the compression tree into a 4-level tree, while MA’s is a 7-level tree. No natural evidence indicates that the assembly index (or MA) corresponds better to how nature works. However, the assembly index is identical to LZ compression [7]. In both cases, the resulting tree of this word problem characterises the same token and is able to reconstruct it in full, without any loss of information, by exploiting redundancy (identical copies) producing a set of possible cause-and-effect chains for which no empirical evidence exists in support of MA. Both LZ and Huffman, just as MA, converge to the same Shannon Entropy rate and can be used to guide a search in chemical space.
Figure 3 shows an illustration of the standard operation of Huffman coding in a typical example, compared to the principle advanced by the AT authors [34]. Proposed in the 50s, the Huffman coding exploits block redundancy by parsing objects, counting block recurrence and was one of the first, if not the first, dictionary-based coding algorithm [28].
As shown in Fig. 3 featuring the ABRACADABRA example, to the left (1A), we see the reconstruction of the sequence from a root node by the method proposed by AT, in general, and in this particular molecular application, when represented as a tree search diagram following binary branching rules. A bifurcation to the right denoted by 1 indicates a new assembly step, whereas a bifurcation to the left indicated by 0 from a node represents a fixed structure (block). The MA algorithm requires seven assembly steps to derive the sequence of interest. However, as shown to the left in Fig. 3, the Huffman coding tree optimises the sequence reconstruction by principles of recursivity in its search compression, as evidenced by the nested bifurcations.
Unlike algorithms like Huffman, MA lacks bifurcations in the assembly search and instead considers a combinatorial search space with a linear sequence progression that cannot be justified by causal chain progression because it may not have anything to do with how an object may have been assembled through many parallel processes. The authors, therefore, conflate causation with plausible assembly ordering.
Hence, it shirks the quantification of emergent hierarchical or nested structures (i.e., modularity optimisation) and intermediate structures within the sequence decomposition/compression. In contrast to MA, the recursiveness observed in complex molecules and biosignatures is detected by RLE and Huffman coding, and it does this in the most optimal way by providing the shortest tree algorithm (what the call ‘assembly pathways’) needed.
The results show that MA performs as any other coding/compression scheme because it is a compression scheme (even if they may have not intended to be so). The predictive power of MA, even if not unique or different, is therefore due to its information-theoretic properties based upon statistical compression of repeated patterns.
Appendix C Methods
The list of MA values for all mass spectral signatures is available in [34]. Various complexity measures were used to classify living versus non-living molecules from the chemical space data in a four-category scheme: natural compounds, metabolites, pharmaceuticals, and industrial compounds, where the natural compounds include the amino acids, consistent with the classifications of Figure 2 in [34]. Further, the mass spectrometry (MS) data of the mixtures (biological extracts and non-living molecules) were categorized into three categories: abiotic, dead, and biological, as consistent with Figure 4 from [34]. The results were subjected to statistical analyses such as the Kolmogorov-Smirnov test, one sample t-tests, and Pearson correlation analysis using GraphPad Prism v. 8.4.3.
Mass spectrometry (MS) data of the extracts and the various molecules used to construct the chemical space for validating MA theory, were analyzed using various complexity measures, including the 1D-string and 2D-matrix Block Decomposition Method (BDM) [55, 56, 58], Shannon’s entropy, and compression algorithms, including Lempel-Ziv-Welch (LZW), All data were first binarised using the online text-to-binary converter with ASCII / UTF-8 character encoding. Run Length Encoding (RLE), Huffman coding, and gzip. The InChI strings of the 99 molecules from (MW vs. MS data) of Fig. 2B, and the 114 molecules from Figure 3 (MS data standard curve) in [34] were binarised and analysed using the OACC (Online Algorithmic Complexity Calculator) app in R, which computed the 1D-BDM (block size of 2, alphabet size of 2, block overlap of zero) and Shannon Entropy scores. The LZW compression lengths were computed with an online LZW calculator using UTF-8 encoding for the 1D strings. Likewise, RLE and Huffman coding compression lengths were obtained using online calculators as additional lossless compression measures to assess the MS bio-signatures. The RLE calculator was set to character then count settings, while the Huffman coding calculator output was set to compression ratio. As for the Figure 2, biological extracts (mixtures) analysis, we used the mass spectra peak matrices of the mixtures (MS2 peaks vs. the total number of peaks) for the above-discussed method/analysis, post-binarisation above the threshold.
For the unpaired (two-samples/independent measures) t-test with Welch’s correction, at a degree of freedom (df) of 100, a critical t-value of 3.390 is expected for a two-tail P-value of 0.001 (i.e., 99.9% confidence). The t-value closest to 3.39 was found for the 1D-BDM and 2D-BDM [56, 53], with a t-value of 6.410 and 6.561, respectively (P $<$ 0.0001), both within the critical region of statistical significance. All complexity measures obtained a non-parametric Kolmogorov-Smirnov test value of P $<$ 0.0001; the Kolmogorov-Smirnov distance D was smallest for the 1D-BDM and 2D-BDM, with both returning a value of 0.707.
Through these statistical assessments, the 1D-BDM and 2D-BDM at a binary conversion threshold of 3 were found to be robust discriminants of molecular complexity in classifying living vs. non-living molecules. The result is shown in Figure 1.
To perform the 2D-BDM on the MS signatures (molecules), the structural distance matrix was extracted from the 2D-molecular structure SDF files for each molecule using the PubChem database. Binary conversion was performed on the matrices after flattening in R at five different conversion thresholds (i.e., -1, 0, 1, 3 and 5). The matrices were flattened by taking rows of values for the distance matrices, or pairs of values for the MS2 matrices (for Fig. 2) and ordering them by rows onto the Rapidtables ASCII/Unicode text string conversion calculator, and removing the spaces. The matrices can be simply pasted as a text, and converted to binary strings using the ASCII/UTF-8 character encoding and Space Output delimiter string settings. The binarised molecular distance matrices were processed by the PyBDM code to obtain the 2D-BDM scores for each molecule. Distance matrices at a binary conversion threshold of 3 were found to be optimal in differential analysis of the MA chemical space signatures and MS signatures into life vs. non-life categories. The matrices at a conversion threshold of 3 were used to compute the 2D-Huffman code and 2D-RLE compression lengths.
C.1 Description of Algorithmic Complexity Measures
The paper by Marshall et al. [32] introduced the molecular assembly (MA) algorithm as a new approach to assess the complexity of biochemical interactomes and distinguish biosignatures of life from abiotic chemistry. The hypothesis is that the molecular assembly (MA) algorithm models the assembly process of biomolecules.
To evaluate the MA algorithm’s utility for analyzing complex biochemical systems, we benchmarked it against established methods from algorithmic information theory that quantify algorithmic complexity. Algorithmic information theory, rooted in computational complexity theory and Kolmogorov complexity, provides a rigorous mathematical framework to assess the information content of objects independent of the specific encoding. We applied several algorithmic information theory techniques to the same datasets analyzed in [32], including molecular distance matrices and mass spectrometry data of chemical interactomes. Our toolkit consisted of:
1D-BDM: The block decomposition method (BDM) segments data vectors into variable blocks and calculates the algorithmic complexity from the entropy of block sizes [56]. 1D-BDM analyzes one-dimensional data such as DNA sequences. 2D-BDM: The two-dimensional BDM acts on distance matrices and other 2D datasets. It divides matrices into rectangular blocks and computes complexity from block entropies.
LZW: The Lempel-Ziv-Welch (LZW) algorithm evaluates complexity by the number of steps needed to compress data (Ziv and Lempel, 1978). It builds a dictionary of sequences and measures complexity by dictionary size. Shannon Entropy: As a measure of uncertainty, or randomness, it quantifies information content based on symbol frequencies in data. Lower entropy indicates simpler patterns.
[…]